# FIVB ranking: Misstep in the right direction

### v1

@@ 2.3 Data
To evaluate the FIVB ranking algorithm, we use the games played by the men’s national teams and published on the FIVB website FIVB (2024). We only consider the games since January 1, 2021, as, before that date, due to the Covid-19 pandemic, many games were eliminated and some had the weighting $\xi_{v}$ incompatible with the official description. [^3] For example, the games played in January 2020 used the weight $\xi_{v}=2.0$ which is not specified in the ranking. The results were collected till the end of December 2023 (with no games after October 2023).
Since the FIVB website publishes the result of the game and the change in the value of skills, $\theta_{m,t+1}-\theta_{m,t}$, we can find the value of $\xi_{v}$ and infer the category of the game from Table 1. To establish the venue of the game, we relied on Wikipedia pages describing the international volleyball events, and matched them with the games from the FIVB website.
@@
  • The games in which, the FIVB ranking displays increments with small absolute value $|\theta_{m,t+1}-\theta_{m,t}|\in\{0,0.01\}$. [^4] There were $67$ games with the absolute value of the increment equal to $0.01$, and $33$ with the increment equal to $0.0$. For example, the Oct. 8, 2023 game China-Poland had the increment equal to $0.01$, and the 24 Sept., 2023, USA-Canada game had the increment equal to $0.0$. This may happen, e.g., when players could not obtain visas, and sometimes it is explained on the FIVB website (e.g., in the case of the Iranian team playing in the USA in 2023) but, in other cases, the non-standard values of $\xi$ are left unexplained. Since these games do not contribute to the change in the rating, excluding them from considerations is consistent with the spirit of the FIVB ranking.
  • The forfeited games we could identify, [^5] Including: i) Denmark’s games in Jan. 2021, ii) Uzbekistan’s and Pakistan’s games in July 2023, and iii) Mongolia’s games in Aug. 2023. even if the FIVB treats them as a “$3-0$” win: because the forfeited games do not reflect on the strength of the team, and thus we do no use them as information for rating.
Thus, we have a total of $M=102$ teams and $T=1151$ games. To characterize the results, we count games with result $y_{t}=y$, played on the neutral- and home-venues
@@
Note that in neutral-venue games, there is no distinction between the home/away teams, so the number of results “$k-l$” (denoted as $y$) and “$l-k$” (denoted as $L-1-y$) should be equal: this is why we count all these results in (29). Although this produces a fractional value $k^{{\textnormal{ntr}}}_{y}$, this formalism simplifies the notation, and we guarantee that $k^{{\textnormal{ntr}}}_{y}=k^{{\textnormal{ntr}}}_{L-1-y}$.
@@
Table 3: Numbers of the FIVB games with outcomes $y\in\mathcal{Y}$, played on the neutral- and home- venues, and their totals.
@@ 2.4 Implicit loss function
˙$\dot{\ell}_{y}(z)$ $=-\frac{\mathcal{N}(z+c_{y})-\mathcal{N}(z+c_{y-1})}{\Phi(z+c_{y})-\Phi(z+c_{y-1})}.$ (32)
@@
To understand what problem it does solve, we note that, since the FIVB still has the structure of the SG optimization, we may treat the function $g^{{\textnormal{FIVB}}}_{y}(z)$ used by (25) as a derivative of an implicit (that is, not explicitly defined) loss function $\ell^{{\textnormal{FIVB}}}_{y}(z)$, i.e., $g^{{\textnormal{FIVB}}}_{y}(z)=\dot{\ell}^{{\textnormal{FIVB}}}_{y}(z)$, and, the latter can be unveiled through the integration of $g^{{\textnormal{FIVB}}}_{y}(z)$, i.e.,
@@
The first is of numerical nature, because, for large $|z|$, both the numerator and the denominator in (32) tend to zero which requires a careful implementation. [^6] Although large values $|z|$ are unlikely to appear in real-world ranking implementation, they do appear when solving optimization (12). In any case, it is preferable to use a numerically stable formulation. We note that even the implementation of (31) is difficult because, in floating-point implementation, for large $z>0$, $\Phi(z+c_{y})$ and $\Phi(z+c_{y-1})$ tend to $1$ and the precision of their difference is constrained by the number of bits in the mantissa representation. We may then exploit the symmetry of the CDF to calculate it only for negative arguments, i.e., $\Phi(z+c_{y})-\Phi(z+c_{y-1})=\Phi(-z-c_{y-1})-\Phi(-z-c_{y})$. However, even with this trick, with growing $|z|$, we need to use the log-domain implementation $\log\big{(}\Phi(z+c_{y})-\Phi(z+c_{y-1})\big{)}$ $=\Psi(z+c_{y})+\log\Big{[}1-\mathrm{e}^{\Psi(z+c_{y-1})-\Psi(z+c_{y})}\Big{]}$ (36) where $\Psi(z)=\log\Phi(z)$, available in numerical software packages, is calculated without explicit evaluation of the CDF. This is merely done to make the calculation feasible. However, the resulting formulas are too complex to produce a practical ranking algorithm.
The second problem is that (32) has a relatively complicated form without obvious interpretation. In that regard, (26) has the advantage: the skills are updated by the difference between the observed and expected numerical scores, and we note that this is also the usual interpretation of the well-known Elo algorithm Elo (2008).
We therefore conjecture that the FIVB rating algorithm was designed to be simple and understandable. The potential drawback is the sub-optimality of the rating results, which we will evaluate in this work.
@@
The second derivative of the implicit loss $\ell^{{\textnormal{FIVB}}}_{y}(z)$ is positive if the first derivative $g^{{\textnormal{FIVB}}}_{y}(z)=\check{r}(z)-r^{\textnormal{FIVB}}_{y}$ increases monotonically in $z$. The latter holds because (28) is monotonically increasing if $r^{\textnormal{FIVB}}_{y+1}<r^{\textnormal{FIVB}}_{y}$, which is true from Table 2.
We note that condition $r^{\textnormal{FIVB}}_{y+1}<r^{\textnormal{FIVB}}_{y}$ is sufficient but not necessary, i.e., it can be violated, and yet we can still obtain a convex function $\ell^{{\textnormal{FIVB}}}_{y}(z)$. The fact that this is possible only reinforces the idea that numerical scores are auxiliary parameters and should not be thought of as a value of the game outcome (because the latter, simply, do not exist).
@@ 3.1 Model identification via cross-validation
The inference (12) may be done if we define (the form of) the loss and the regularization functions, as well as, if we find their parameters $\boldsymbol{p}$ (also called hyper-parameters, in the machine learning language) which affect all the functions describing the models, i.e., we can write $\ell^{{\textnormal{loss}}}(z)\equiv\ell^{{\textnormal{loss}}}_{y}(z;\boldsymbol{p})$, $\ell_{y}(z)\equiv\ell_{y}(z;\boldsymbol{p})$ and $\rho(\boldsymbol{\theta})\equiv\rho(\boldsymbol{\theta};\boldsymbol{p})$.
A well-known approach to finding the hyper-parameters, particularly suitable for relatively small data sets, relies on cross-validation, where we split the data into multiple training and validation sets (Barber, 2012, Ch. 13.2.2), and we verify how well the results (12) match the validation data using a metric $\ell^{{\textnormal{val}}}_{y}(\boldsymbol{x}^{\mathsf{T}}\hat{\boldsymbol{\theta}})$.
If we know the model (1), and especially if we use the log-score as the loss function, i.e., in $\ell^{{\textnormal{loss}}}_{y}(z)=\ell_{y}(z)$, the validation metric used in practice is also the log-score
$\ell^{{\textnormal{val}}}_{y}(z)$ $=\ell_{y}(z);$ (37)
this is the metric we use here as well.
We will use the leave-one-out (LOO) cross-validation, where the validation sets contain only one sample and thus, we can consider all possible $T$ validation/testing splits, (Hastie et al., 2009, Ch. 2.9)
@@
$\hat{\boldsymbol{\theta}}_{\backslash{\tau}}$ $=\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}\big{[}J(\boldsymbol{\theta})-\ell^{{\textnormal{loss}}}_{y_{\tau}}(\boldsymbol{x}^{\mathsf{T}}_{\tau}\boldsymbol{\theta};\boldsymbol{p})\big{]}.$ (41)
@@
which is the geometric mean of the probabilities of observing $Y_{t}=y_{t}$, calculated from $\hat{\boldsymbol{\theta}}_{\backslash{t}}$. Thus, $V(\boldsymbol{p})$ is potentially easier to interpret than $U(\boldsymbol{p})$; and, with linearization, for small $U(\boldsymbol{p})$, e.g., $U(\boldsymbol{p})<0.1$, we have $V(\boldsymbol{p})\approx 1-U(\boldsymbol{p})$.
Although using the LOO strategy eliminates the ambiguity of how the validation sets are selected (we test all, $T$ possible testing/validations splits), solving $T$ times the optimization problem (41) becomes computationally demanding.
Thus, instead, we use the approximate leave-one-out (ALO) approach (Beirami et al., 2017, Rad and Maleki, 2020, Burn, 2020): we solve (12) once to find $\hat{\boldsymbol{\theta}}\equiv\hat{\boldsymbol{\theta}}(\boldsymbol{p})$ (this is where most of the computational complexity lies), and make a quadratic approximation of the function $J(\boldsymbol{\theta})$ around $\hat{\boldsymbol{\theta}}$, which allows us to find the closed-form approximation of (40) as follows:
@@
We may use, of course, $\ell_{y_{t},v_{t}}^{{\textnormal{loss}}}(z_{t};\boldsymbol{p})$ or $\ell_{y_{t},h_{t}}^{{\textnormal{loss}}}(z_{t};\boldsymbol{p})$ defined in (18) or (19), and then the first/second derivatives used in (43) should be calculated as
˙$\dot{\ell}_{y_{t},v_{t}}^{{\textnormal{loss}}}(z_{t};\boldsymbol{p})$ $=\xi_{v_{t}}\dot{\ell}_{y_{t}}^{{\textnormal{loss}}}(z_{t};\boldsymbol{p}),$ (44)
¨$\ddot{\ell}_{y_{t},v_{t}}^{{\textnormal{loss}}}(z_{t};\boldsymbol{p})$ $=\xi_{v_{t}}\ddot{\ell}_{y_{t}}^{{\textnormal{loss}}}(z_{t};\boldsymbol{p}),$ (45)
˙$\dot{\ell}_{y_{t},h_{t}}^{{\textnormal{loss}}}(z_{t};\boldsymbol{p})$ =˙$=\dot{\ell}_{y_{t}}^{{\textnormal{loss}}}(z_{t}+h_{t}\eta;\boldsymbol{p}),$ (46)
¨$\ddot{\ell}_{y_{t},h_{t}}^{{\textnormal{loss}}}(z_{t};\boldsymbol{p})$ =¨$=\ddot{\ell}_{y_{t}}^{{\textnormal{loss}}}(z_{t}+h_{t}\eta;\boldsymbol{p}).$ (47)
If we predefine candidate hyperparameters in the set $\mathcal{P}=\{\boldsymbol{p}_{{\textnormal{a}}},\boldsymbol{p}_{{\textnormal{b}}},\boldsymbol{p}_{{\textnormal{c}}},\ldots\}$, e.g., using those specified in the existing ranking, or modifying them using a particular insight into the ranking problem, the optimization is simple:
$\hat{\boldsymbol{p}}=\mathop{\mathrm{argmin}}_{\boldsymbol{p}\in\mathcal{P}}U(\boldsymbol{p})$ (48)
and is essentially done by comparing the value of $U(\boldsymbol{p})$ for all candidates in $\mathcal{P}$.
On the other hand, optimizing the value of the hyperparameters $\boldsymbol{p}$ as suggested by (38) is more difficult but can be done with appropriate numerical tools, as we explain in Appendix B.
Empirical entropy as a comparison baseline
Model identification results can be compared to a simple baseline, where we assume that the prediction of the game result $Y_{t}$ is based on the empirical frequency of the results observed in all other games, i.e.,
~$\tilde{\ell}^{{\textnormal{val}}}_{y_{t}}=-\log\frac{k_{y_{t}}-1}{T-1},$ (49)
where
$k_{l}=k^{{\textnormal{ntr}}}_{l}+k^{{\textnormal{hfa}}}_{l},$ (50)
is the venue-independent game count. [^7] In (49), $k_{l}$ and $T$ must be decreased by one, because, making probabilistic prediction of the outcome $Y_{t}$, its actual result, $y_{t}$, should be assumed unknown when calculating the empirical frequency.
Then, the frequency-based performance baseline (39) is found as
$\tilde{U}$ $=-\frac{1}{T}\sum_{t=1}^{T}\log\frac{k_{y_{t}}-1}{T-1}$ (51)
$=-\sum_{l=0}^{L-1}\frac{k_{l}}{T}\log\frac{k_{l}-1}{T-1}=H(\boldsymbol{q},T)$ (52)
where $\boldsymbol{q}=[q_{0},\ldots,q_{L-1}]$, $q_{l}=k_{l}/T$ is the empirical frequency of events $Y_{t}=l$, and
$H(\boldsymbol{q},T)=-\sum_{l=0}^{L-1}q_{l}\log\frac{q_{l}-1/T}{1-1/T};$ (53)
for large $T$, (53) tends to the entropy $H(\boldsymbol{q})$ of the source which generates the outcomes $Y_{t}$ with probabilities defined in the vector $\boldsymbol{q}$.
If we take into account the HFA, (51) becomes
$\tilde{U}$ $=\frac{1}{T}\Big{[}T^{{\textnormal{ntr}}}\tilde{U}^{{\textnormal{ntr}}}+T^{{\textnormal{hfa}}}\tilde{U}^{{\textnormal{hfa}}}\Big{]}$ (54)
$\tilde{U}^{{\textnormal{ntr}}}$ $=H(\boldsymbol{q}^{{\textnormal{ntr}}},T^{{\textnormal{ntr}}})$ (55)
$\tilde{U}^{{\textnormal{ntr}}}$ $=H(\boldsymbol{q}^{{\textnormal{hfa}}},T^{{\textnormal{hfa}}})$ (56)
where, $T^{{\textnormal{ntr}}}=\sum_{l=0}^{L-1}k^{{\textnormal{ntr}}}_{l}$, $T^{{\textnormal{hfa}}}=\sum_{l=0}^{L-1}k^{{\textnormal{hfa}}}_{l}$, $\boldsymbol{q}^{{\textnormal{ntr}}}=[k^{{\textnormal{ntr}}}_{0}/T^{{\textnormal{ntr}}},\ldots,k^{{\textnormal{ntr}}}_{L-1}/T^{{\textnormal{ntr}}}]$, and $\boldsymbol{q}^{{\textnormal{hfa}}}=[k^{{\textnormal{hfa}}}_{0}/T^{{\textnormal{hfa}}},\ldots,k^{{\textnormal{hfa}}}_{L-1}/T^{{\textnormal{hfa}}}]$.
These entropy-like metrics, shown in Table 4, are performance upper bounds in the sense that any well-designed rating strategy, which uses $\hat{\boldsymbol{\theta}}_{\backslash{t}}$ as a basis for predicting the outcomes, should provide the cross-validation metric $U(\boldsymbol{p})$ that is smaller than the value $\tilde{U}$ shown in Table 4a, and larger than the corresponding average probabilities $\tilde{V}$ shown in Table 4b.
$\tilde{U}$ $\tilde{U}^{{\textnormal{ntr}}}$ $\tilde{U}^{{\textnormal{hfa}}}$
1.68 1.69 1.66
a)
$\tilde{V}$ $\tilde{V}^{{\textnormal{ntr}}}$ $\tilde{V}^{{\textnormal{hfa}}}$ 0.19 0.18 0.19
b)
Table 4: Empirical entropy defined in (54), (55), and (56), calculated for the games FIVB, as well as their transformation via (42).
Similarly, the reference metrics $\tilde{U}^{{\textnormal{ntr}}}$ and $\tilde{U}^{{\textnormal{hfa}}}$ should be treated as trivial upper bounds on the cross-validation results conditioned on the game venue
Of course, we can recover (39) as $U(\boldsymbol{p})=[T^{{\textnormal{ntr}}}U^{{\textnormal{ntr}}}(\boldsymbol{p})+T^{{\textnormal{hfa}}}U^{{\textnormal{hfa}}}(\boldsymbol{p})]/T$, see also (54).

### v2

@@ 2.4 Data
To evaluate the FIVB ranking algorithm, we use the matches played by the men’s national teams and published on the FIVB website (FIVB, 2024). We only consider the matches since January 1, 2021, as, before that date, due to the Covid-19 pandemic, many matches were eliminated and some had the weighting $\xi_{v}$ incompatible with the official description. [^5] For example, the matches played in January 2020 used the weight $\xi_{v}=2.50$ which is not specified in the ranking. The results were collected till the end of December 2023 (with no matches after October 2023).
Since the FIVB website publishes the result $y_{t}$ of the match $t$ and the change in the value of skills, $\theta_{m,t+1}-\theta_{m,t}$, we can find the value of $\xi_{v_{t}}$ and infer the match category $v_{t}$ from Table 1. To establish the venue of the match (which affects $h_{t}$), we relied on Wikipedia pages describing the international volleyball events that we paired with matches from the FIVB website. [^6] These results were non exhaustively validated by comparing with the venues of main international events shown on the FIVB website (such as Nations League, Challenger Cup or World Championship).
@@
  • The matches in which, the FIVB ranking displays increments with small absolute value $|\theta_{m,t+1}-\theta_{m,t}|\in\{0,0.01\}$. [^7] There were $67$ matches with the absolute value of the increment equal to $0.01$, and $33$ with the increment equal to $0.0$. For example, the Oct. 8, 2023 match China-Poland had the increment equal to $0.01$, and the 24 Sept., 2023, USA-Canada match had the increment equal to $0.0$. This may happen, e.g., when players could not obtain visas, and sometimes it is explained on the FIVB website (e.g., in the case of the Iranian team playing in the USA in 2023) but, in other cases, the non-standard values of $\xi$ are left unexplained. Since these matches do not contribute to the change in the ranking, excluding them from considerations is consistent with the spirit of the FIVB ranking.
  • The forfeited matches we could identify, [^8] Including: i) Denmark’s matches in Jan. 2021, ii) Uzbekistan’s and Pakistan’s matches in July 2023, and iii) Mongolia’s matches in Aug. 2023. Even if FIVB treats a forfeited match as a “$3-0$” win, in our opinion, the match which did not take place does not reflect on the strength of the team, and thus we do not use it as information for ranking.
Then, we have a total of $M=102$ teams and $T=1151$ matches. To characterize the results, we count matches with result $y_{t}=y$, played on the neutral- and home-venues
@@
Note that in neutral-venue matches, there is no distinction between the home/away teams, so the number of results “$k-l$” (denoted as $y$) and “$l-k$” (denoted as $L-1-y$) should be equal: this is why we count all these results in (24). Although this produces a fractional value $k^{{\textnormal{ntr}}}_{y}$, this formalism simplifies the notation, and we guarantee that $k^{{\textnormal{ntr}}}_{y}=k^{{\textnormal{ntr}}}_{L-1-y}$.
@@
Table 3: Numbers of the FIVB matches with outcomes $y\in\mathcal{Y}$, played on the neutral- and home- venues, and their totals.
@@ 2.5 Implicit loss function
˙$\dot{\ell}_{y}(z)$ $=-\frac{\mathcal{N}(z+c_{y})-\mathcal{N}(z+c_{y-1})}{\Phi(z+c_{y})-\Phi(z+c_{y-1})},$ (27)
where we use the Gaussian probability density function (PDF)
$\mathcal{N}(z)=\dot{\Phi}(z)=\frac{1}{\sqrt{2\pi}}\exp(-0.5z^{2}).$ (28)
@@
To understand what problem it does solve, we note that, since the FIVB ranking (20) has the structure of the SG optimization (12), we may treat the function $g^{{\textnormal{FIVB}}}_{y}(z)$ used by (20) as a derivative of an implicit (that is, not explicitly defined) loss function $\ell^{{\textnormal{FIVB}}}_{y}(z)$, i.e., $g^{{\textnormal{FIVB}}}_{y}(z)=\dot{\ell}^{{\textnormal{FIVB}}}_{y}(z)$, and, the latter can be unveiled through the integration of $g^{{\textnormal{FIVB}}}_{y}(z)$, i.e.,
@@
The first is of numerical nature, because, for large $|z|$, both the numerator and the denominator in (27) tend to zero which requires a careful implementation.
The second problem is that (27) has a relatively complicated form without obvious interpretation. In that regard, (21) has a practical advantage: the skills $\boldsymbol{\theta}_{t}$ are updated using the difference between the observed and expected numerical scores. Note that this is also the usual interpretation of the well-known Elo algorithm (Elo, 2008).
We therefore conjecture that the FIVB ranking algorithm was designed to be simple and understandable, which is a typical requirement in the sport ranking (Csató, 2024, Sec. 1). The potential drawback is the sub-optimality of the ranking results, which we will evaluate in this work.
@@
The second derivative of implicit loss $\ell^{{\textnormal{FIVB}}}_{y}(z)$ is positive if the first derivative $g^{{\textnormal{FIVB}}}_{y}(z)=\check{r}(z)-r^{\textnormal{FIVB}}_{y}$ increases monotonically in $z$. The latter holds because (23) is monotonically increasing if $r^{\textnormal{FIVB}}_{y+1}<r^{\textnormal{FIVB}}_{y}$, which is true from Table 2.
We note that condition $r^{\textnormal{FIVB}}_{y+1}<r^{\textnormal{FIVB}}_{y}$ is sufficient but not necessary, i.e., it can be violated, and yet we can still obtain a convex function $\ell^{{\textnormal{FIVB}}}_{y}(z)$ (see examples in Sec. 3.3.2). The fact that this is possible only reinforces the idea that numerical scores are auxiliary parameters and should not be thought of as a value of the match outcome (because the latter simply do not exist).
@@ 3.1 Model identification via cross-validation
Inference (7) may be done if we define (the form of) the loss and regularization functions, as well as, if we find their parameters $\boldsymbol{p}$ (also called hyper-parameters, in the machine learning language) that affect all the functions describing the models, i.e., we can write $\ell^{{\textnormal{loss}}}(z)\equiv\ell^{{\textnormal{loss}}}_{y}(z;\boldsymbol{p})$, $\ell_{y}(z)\equiv\ell_{y}(z;\boldsymbol{p})$ and $\rho(\boldsymbol{\theta})\equiv\rho(\boldsymbol{\theta};\boldsymbol{p})$.
A well-known approach to finding the hyper-parameters, particularly suitable for relatively small data sets, relies on leave-one-out (LOO) cross-validation (Hastie et al., 2009, Ch. 2.9), where we remove one observation $y_{t}$, and we verify how well the results (7), denoted as $\hat{\boldsymbol{\theta}}_{\backslash{t}}$, match the removed data using a metric $\ell^{{\textnormal{val}}}_{y_{t}}(\boldsymbol{x}_{t}^{\mathsf{T}}\hat{\boldsymbol{\theta}}_{\backslash{t}})$; for the latter, most often we use the log-score $\ell^{{\textnormal{val}}}_{y}(z)=\ell_{y}(z)$. This is repeated for all $T$ samples and may be defined as follows (Hastie et al., 2009, Ch. 2.9)
@@
$\hat{\boldsymbol{\theta}}_{\backslash{t}}$ $=\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}\big{[}J(\boldsymbol{\theta})-\ell^{{\textnormal{loss}}}_{y_{t}}(\boldsymbol{x}^{\mathsf{T}}_{t}\boldsymbol{\theta};\boldsymbol{p})\big{]}.$ (35)
where $U(\boldsymbol{p})$ is the averaged validation metric. [^9] Note that, while the LOO cross-validation uses the validation sets containing only one element, in general, we can use larger sets. However, they must then be defined arbitrarily (e.g., randomly), as it is rarely possible to enumerate all of them. On the other hand, in the LOO approach we do enumerate all $T$ validation sets. Thus, given the data, the results are independent of the random/arbitrary definition of the validation sets. This removes any ambiguity in the numerical optimization of $U(\boldsymbol{p})$ required in this work.
We can also calculate the metrics for the matches played on the neutral and the home venues, given, respectively, by
where, $U(\boldsymbol{p})=\big{(}T^{{\textnormal{ntr}}}U^{{\textnormal{ntr}}}(\boldsymbol{p})+T^{{\textnormal{hfa}}}U^{{\textnormal{hfa}}}(\boldsymbol{p})\big{)}/T$.
@@
which is the geometric mean of the probabilities of observing $Y_{t}=y_{t}$, calculated from $\hat{\boldsymbol{\theta}}_{\backslash{t}}$. Thus, $V(\boldsymbol{p})$ is potentially easier to interpret than $U(\boldsymbol{p})$; and, with linearization, for small $U(\boldsymbol{p})$, e.g., $U(\boldsymbol{p})<0.1$, we have $V(\boldsymbol{p})\approx 1-U(\boldsymbol{p})$. For example, $U(\boldsymbol{p})=1.4$ yields $V(\boldsymbol{p})\approx 24.7$%, while for a uniform distribution $\mathsf{P}_{y}(z)=\frac{1}{6},\forall y$ (i.e., $V(\boldsymbol{p})=16.7$%) we have $U(\boldsymbol{p})=1.79$. The values $V(\boldsymbol{p})$ are shown on the right-hand auxiliary axis in Figs. 3, Fig. 5, and Fig. 7.
While the complexity is reduced, we still need to solve (35) $T$ times, and, to alleviate the complexity, we apply here the approximate leave-one-out (ALO) approach (Beirami et al., 2017),(Rad and Maleki, 2020), (Burn, 2020), where (7) is solved once to find $\hat{\boldsymbol{\theta}}\equiv\hat{\boldsymbol{\theta}}(\boldsymbol{p})$ (this is where most of the computational complexity lies), and we make a quadratic approximation of the function $J(\boldsymbol{\theta})$ around $\hat{\boldsymbol{\theta}}$, which allows us to find the closed-form approximation of (34) as follows: